A bar 2 m long makes an angle of 30° with the horizontal. A force of 40 is applied 0.4 m from the upper end. Calculate the torque due to this a force about each end.


1
Expert's answer
2021-05-09T13:06:54-0400
"\tau_{upper}=(40)(0.4)\cos{30}=13.9\ Nm\\
\tau_{lower}=(40)(1.6)\cos{30}=55.4\ Nm"